OpenStudy (anonymous):

Find the distance between the lines with equations -4x+7y+9=0 and y=4/7-9

OpenStudy (anonymous):

I THINK this is the formula , but i'm not sure, @ganeshie8 tell me if i'm wrong or not... \[d = \frac{ |b _{2}-b _{1}| }{ \sqrt{m ^{2}+1} }\]

OpenStudy (anonymous):

and this formula is for common form for standard form:

OpenStudy (anonymous):

\[d = \frac{ |c _{2}-c _{1}| }{ \sqrt{a ^{2}+a ^{1}} }\]

OpenStudy (anonymous):

i dont understand all this caculator stuff

OpenStudy (anonymous):

and you should know that these formulas are used for \( \Huge {Straight \; Lines} \)

OpenStudy (anonymous):

ok,meybe i can't explain that, @DebbieG , @ganeshie8

OpenStudy (anonymous):

lol

OpenStudy (debbieg):

If they are parallel lines, the distance between them is a constant. I would think the easiest way to find that distance would be to find the difference between the y-intercepts.

OpenStudy (anonymous):

screw it. i already submitted. thanks for trying